### v2

# MTL-KD: Multi-Task Learning Via Knowledge Distillation for Generalizable Neural Vehicle Routing Solver

Abstract
Multi-Task Learning (MTL) in Neural Combinatorial Optimization (NCO) is a promising approach to train a unified model capable of solving multiple Vehicle Routing Problem (VRP) variants. However, existing Reinforcement Learning (RL)-based multi-task methods can only train light decoder models on small-scale problems, exhibiting limited generalization ability when solving large-scale problems. To overcome this limitation, this work introduces a novel multi-task learning method driven by knowledge distillation (MTL-KD), which enables the efficient training of heavy decoder models with strong generalization ability. The proposed MTL-KD method transfers policy knowledge from multiple distinct RL-based single-task models to a single heavy decoder model, facilitating label-free training and effectively improving the model’s generalization ability across diverse tasks. In addition, we introduce a flexible inference strategy termed Random Reordering Re-Construction (R3C), which is specifically adapted for diverse VRP tasks and further boosts the performance of the multi-task model. Experimental results on 6 seen and 10 unseen VRP variants with up to 1000 nodes indicate that our proposed method consistently achieves superior performance on both uniform and real-world benchmarks, demonstrating robust generalization abilities.
1 Introduction
The Vehicle Routing Problem stands as a recognized classical challenge within the field of combinatorial optimization. Due to its widespread applicability across numerous real-world scenarios [1, 2, 3], including logistics distribution, traffic scheduling, and emergency response, the VRP has consistently been a subject of intense research interest. However, traditional optimization algorithms [4, 5, 6] often necessitate intricate domain-specific knowledge for their design and exhibit limitations in efficiency when applied to large-scale instances, thereby hindering their scalability in practical applications.
In recent years, Neural Solvers [7, 8, 9, 10] have demonstrated significant potential in tackling VRPs. These models offer high computational efficiency and reduced reliance on domain-specific knowledge, achieving competitive performance with traditional algorithms. However, the inherent design of these single-task models significantly restricts their versatility in handling diverse problem instances. The necessity to retrain the model even for slightly different problem settings severely limits their practical utility.
The diverse nature of VRPs in the real world has spurred recent research efforts to develop unified models capable of solving multiple VRP variants and exhibiting good generalization on unseen tasks[11, 12, 13]. Currently, the majority of unified models for VRP variants adopt a heavy encoder-light decoder (HELD) architecture. These models have relatively low computational costs and can be directly trained using RL, demonstrating excellent performance on small-scale problems. However, their generalization ability significantly deteriorates when dealing with large-scale instances, primarily because the light decoder struggles to extract sufficient information from the high-density and complex node embeddings[14]. Furthermore, training them on large-scale instances also presents the challenge of excessively large gradient information, collectively leading to the inadequacy of existing VRP variant models in handling large-scale problems.
On the other hand, the light encoder-heavy decoder[15, 16, 17] architecture has demonstrated excellent generalization capabilities on large-scale problems. Its iterative decoding process, where the heavy decoder continuously re-evaluates the relationships between the remaining nodes, coupled with the dynamic change in the number of nodes, contributes to the strong scale generalization performance of this architecture. To achieve scale generalization within a multi-task unified model, we aim to leverage the heavy decoder architecture. However, the substantial memory and computational demands inherent in heavy decoders render direct RL impractical; they typically employ supervised learning, which necessitates extensive labeled data, a demand further amplified by multi-task training. While SIT[17] employs a self-improved Training method using operations like local reconstruction [15] to refine model-generated solutions as training labels, this suffers from the generation of numerous low-quality labels during early and mid-training, leading to a protracted process and an increased training burden due to time-consuming pseudo-label generation. We provide a comprehensive literature review on autoregressive neural solvers and multi-task neural solvers for VRPs in Appendix A.
To address the limitations of existing supervised training methods for heavy decoders, we introduce knowledge distillation as a novel training paradigm for the heavy decoder model, thereby circumventing the need for extensive label generation. We employ easily RL-trained HELD teacher models as supervision to train our multi-task heavy decoder model, thus achieving label-free training for the heavy decoder. The heavy decoder model trained with this method demonstrates strong generalization performance across both scales and tasks. Furthermore, during inference, we propose a general Random Reordering Reconstruction strategy for various VRP variants. By randomly reordering the external order of subtours, R3C significantly enhances solution sampling diversity, mitigates the risk of getting trapped in local optima, and further improves performance. Our contributions are summarized as follows: 1) We achieve efficient label-free training of heavy decoder models for multi-task VRPs through knowledge distillation. 2) We propose a novel R3C strategy to further enhance the performance of the multi-task model. 3) The MTL-KD model demonstrates excellent performance on seen training tasks, unseen tasks, and real-world datasets, exhibiting good scale generalization ability and significantly outperforming existing multi-task VRP models.
2 Preliminaries
Capacitated Vehicle Routing Problem and Its Variants
The Capacitated Vehicle Routing Problem (CVRP) is typically defined as follows: Given a central depot $v_{0}$ and $N$ customer nodes $v_{1}$ to $v_{n}$, all interconnected with distances $e_{ij}$. Each customer $i$ has a demand $d_{i}$, and all vehicles depart from the depot, serve customers, and return to the depot. Each vehicle has a capacity $C$, satisfying $C>d_{i}$. All customers must be visited exactly once, and the total load of the vehicle during its route must not exceed its capacity. The objective of CVRP is to minimize the total travel distance of all vehicles.
By introducing additional constraints, CVRP can be extended into various variants. The four types of additional constraints studied in this paper include: (1) Open Route (O): Vehicles are not required to return to the depot after completing their service. (2) Time Windows (TW): Each node $i$ has a service time window $[s_{i},l_{i}]$, within which vehicles must arrive at the customer (early arrivals require waiting until the earliest service start time). (3) Backhaul (B): Customer demands can be positive (linehauls) or negative (backhauls), and there is no restriction on the sequence of visiting linehauls and backhauls customers. (4) Duration Limit (L): The total travel distance of each vehicle must not exceed a predefined threshold $L$. Combining these constraints with CVRP can form 16 different VRP variants, with specific details provided in Appendix B.
Solution Construction Process by Neural Solver
Constructive neural solvers [7, 8, 15] typically employ an encoder-decoder architecture and construct solutions in an autoregressive manner. For a given VRP instance $\mathcal{G}=(V,E)$, where $V$ denotes the set of nodes, including node features such as demand, and $E$ represents the set of edges, often formed by a distance matrix. The encoder is used to extract node information and generate embedding vectors for each node, while the decoder constructs the next node based on the current partial solution state $s_{t-1}$. The decoder is subject to node constraints during the construction process to avoid generating infeasible solutions. The entire solving process can be represented as:
$\pi_{\theta}(\tau|\mathcal{G})=\prod_{t=1}^{\ell}\pi_{\theta}(a_{t}|s_{t},\mathcal{G}),$
where $\tau$ is the complete solution, $a_{t}$ is the action (node) selected at time step $t$, and $\ell$ is the total number of actions. The decoder continuously constructs nodes until all nodes have been visited, forming a complete solution. Through training, the model learns how to effectively select the next node based on the current partial solution, thereby optimizing the quality of the entire solution.
Knowledge distillation
Knowledge Distillation[18] is a model compression and knowledge transfer technique aimed at transferring the knowledge learned by a high-performing, pre-trained teacher model to a smaller student model. The core idea is to achieve this transfer by minimizing the discrepancy between the outputs of the teacher and student models. The student model’s training loss typically includes a distillation loss ($L_{KD}$) that measures the difference between the student’s predictions and the teacher’s soft targets, generally the KL divergence $L_{KD}=\text{KL}(\pi_{T}||\pi_{S})$. Additionally, depending on the specific task requirements, an original task loss ($L_{Task}$) can optionally be included, and the total loss ($L$) for the student model can be expressed as:
$L=\alpha L_{KD}+(1-\alpha)L_{Task}\quad$
where ($\alpha$) is a weight parameter. In the field of Neural Combinatorial Optimization, knowledge distillation has been used to achieve model lightweighting and can improve the generalization ability of models on unseen problems to a certain extent[19, 20, 21, 22].
3 Multi-Task Learning Via Knowledge Distillation
In this section, we propose a multi-task learning framework via knowledge distillation to address the challenge of applying heavy decoder models in the multi-task domain due to their reliance on large amounts of labeled data. The proposed framework effectively enhances the model’s generalization capabilities across different tasks and scales.
3.1 Framework Overview
Our proposed multi-task training framework via knowledge distillation aims to enhance the model’s generalization capabilities across different tasks and scales. As depicted in Figure 1, the framework first categorizes VRP variants into seen and unseen tasks for model training. Given $N$ seen tasks, we first pre-train $N$ individual teacher models employing a Light Encoder-Heavy Decoder architecture[15]. Subsequently, we construct a multi-task heavy decoder student model and train it using knowledge distillation, leveraging the output distributions of the teacher models as supervision signals.
Figure 1: Framework for Multi-Task Training via Knowledge Distillation. Top-left: seen and unseen tasks. Bottom-left: Independent training of teacher models. Right: Student model distillation via teacher output distributions.
3.2 Teacher Model Training
For each seen VRP variant task, we first generate the corresponding instance data. Then, we independently train a teacher model for each task, which adopts the POMO[8] structure and serves as a policy network. We optimize the parameters of the teacher models using the policy gradient method in RL, with the objective of maximizing the reward (negative tour length). The loss function[23] can be expressed as:
$L(\theta^{T})=-\mathbb{E}_{\tau\sim\pi_{\theta^{T}}(\cdot|\mathcal{G})}[(r(\tau)-b^{T}(\mathcal{G}))\log\pi_{\theta^{T}}(\tau|\mathcal{G})]$
where $b^{T}(\mathcal{G})$ is the baseline (average reward of multi-start trajectories for the instance), $\pi_{\theta^{T}}$ is the teacher model’s policy distribution, $\theta^{T}$ represents the parameters, and $r$ is the reward.
3.3 Student Model Training
During the student model training phase, we construct a multi-task heavy decoder model. In each training batch, we simultaneously generate instance data for the $N$ seen problem types and feed them into the student model for processing. At each decoding step $t$, the student model outputs a predicted probability distribution over the next nodes $\pi_{\theta^{S}}(a_{t}|s_{t},\mathcal{G})$, encompassing the probability distributions for the $N$ problem types. Concurrently, we input each problem instance to its corresponding pre-trained teacher model to obtain the teacher model’s probability distribution $\pi_{\theta^{T}n}(a_{t}|s_{t},\mathcal{G})$ under the same state $s_{t}$. We calculate the Kullback-Leibler divergence between the student model’s output probability distribution and the corresponding teacher model’s output probability distribution as the knowledge distillation loss $L_{KD}^{(t)}$:
$L_{KD}^{(t)}=\sum_{n=1}^{N}\text{KL}(\pi_{\theta^{T}n}(a_{t}|s_{t},\mathcal{G})\|\pi_{\theta^{S}}(a_{t}|s_{t},\mathcal{G}))$
By minimizing this loss function, the student model learns to mimic the behavior of multiple teacher models, thereby acquiring more generalized knowledge and improving its performance in multi-task learning and cross-task generalization. After each decoding step, we select the next node based on the student model’s probability distribution, transition to the new state, and repeat the aforementioned alignment process.
4 Architecture of Generalizable Multi-Task Neural Solver
Heavy decoder models have demonstrated remarkable scale generalization performance on single-task VRP. However, existing multi-task models for VRP variants predominantly rely on light encoder-heavy decoder architectures, exhibiting poor scale generalization. To address this limitation, as depicted in Figure 2, we have developed a multi-task heavy decoder model specifically designed for VRP variants, aiming to overcome the inadequate generalization capabilities of current approaches.
4.1 Encoder
Given an instance $S$ of the VRP, which includes $N$ node features $(s_{1}...s_{n})$, in VRP variants, the node features are uniformly represented as $(x,d,\delta,s,l)$, denoting the coordinate information, demand, service time, and the start and end times of the time window, respectively. These node features are then mapped through a linear layer to obtain the initial embedding $h^{0}=(h_{1}^{0}...h_{n}^{0})$. Subsequently, these initial embeddings are passed through a Transformer layer to capture node relationships and generate the node embeddings. Therefore, the output of the encoder can be represented as:
$H=\text{TF}(h^{0})=(h_{1},...,h_{n})$
where $H$ is the matrix of final node embeddings, and $h_{i}$ represents the embedding of the $i$-th node.
4.2 Decoder
During the decoding phase, we first extract the dynamic features $D=\{l,t,d,o\}$, encompassing the current vehicle load, current time, remaining route duration, and a binary flag indicating an open route. These dynamic features are combined with last visited node($h_{last}$) and the depot node($h_{depot}$) respectively, and then each combination is passed through its own linear layer. The resulting embeddings, along with the unvisited nodes($h_{unvisited}$), are processed by an $L$-layer Transformer network to yield updated node embeddings:
$h^{L}=\text{TF}([h_{unvisited},\text{Linear}(D\oplus h_{last}),\text{Linear}(D\oplus h_{depot})])$
Next, we compute compatibility scores between the embeddings of all unvisited nodes and a context embedding $h_{q}=\text{Linear}(\text{concat}(h_{last},h_{depot}))$ using single-head attention:
$c(h_{i}^{L},h_{q})=\text{SHA}(h_{i}^{L},h_{q})$
To ensure valid routes, a mask vector mask is applied to these scores. Finally, the probability $\pi(i|s_{t})$ of selecting the $i$-th unvisited node is determined by the softmax function applied to the masked compatibility scores:
$\pi(i|s_{t})=\text{softmax}(c(h_{i}^{L},h_{q})+\text{mask}_{i})$
where $c(h_{i}^{L},h_{q})$ is the compatibility score, $h_{i}^{L}$ is the embedding of the $i$-th unvisited node from $h^{L}$, and $s_{t}$ is the current decoder state. The mask vector contains 0 for valid unvisited nodes and $\infty$- ∞ for invalid ones.
To address the varying number of unvisited nodes across different instances within a batch, which can impact batch processing efficiency, we utilize padding to unify the sequence length and a masking mechanism to remove the influence of the padded depot node. Additionally, layer normalization [15] has been removed from all attention layers.
Figure 2: Architecture of the proposed multi-task neural solver. Left to right: Encoder, Decoder, Node Padding & Masking.
5 Random Reordering Re-Construct
Random Re-Construct (RRC) [15] is an iterative method to improve solution quality by randomly sampling and re-optimizing segments of the initial solution. RRC enhances sampling diversity via random subtour reversals and segment lengths. However, reversals can violate feasibility for some problems (e.g., VRPTW), and sampling solely on the original sequence limits diversity, hindering iterative performance. We thus propose Random Reordering Re-Construct to increase sampling diversity and improve iterative performance.
The R3C strategy is detailed in Figure 3. Given a solution, we first decompose it into subtours, then randomize their external order in the sequence. Subsequently, a random-length contiguous segment is randomly sampled and re-optimized by our model. Improved segments replace the original. By randomizing subtour order, R3C allows for more random combinations of subtours into partial solutions, aiding escape from local optima and improving iterative performance. All sampled segments end at the depot. Additionally, feasible subtour reversals are also incorporated to further enhance search diversity.
Figure 3: Comparison of RRC and R3C methods. RRC increases the diversity of sampled subproblems by randomly reversing subtours, while R3C enhances this diversity by randomly reordering the external sequence of subtours.
6 Experiments
In this section, we present a series of experiments conducted on 16 VRP variants to validate the performance of our model. All experiments are performed using one NVIDIA RTX 4090 GPU.
Baselines
The baseline methods are categorized into two groups: traditional heuristic solvers and multi-task neural solvers. (1) Traditional Solvers: PyVRP [24]: An open-source Hybrid Genetic Search (HGS)-based solver supporting multiple complex VRP variants. It represents the current state-of-the-art in heuristic approaches for the problems considered in this study; OR-Tools [6]:Google’s open-source optimization toolkit, applicable to all VRP variants examined in this study. (2) Multi-task Neural Solvers: MT-POMO [11]: Multi-task extension of the POMO framework, enabling simultaneous learning across different VRP variants; MVMoE[12]: An architecture that enhances model capacity and improves performance over MT-POMO by incorporating Mixture-of-Experts (MoE) modules. Given the differences in problem variants and visible task training among the baseline models, we only compare methods with the same training setup.
Datasets
We adopte the same problem settings as MVMoE[12]. We generate a test set comprising 16 VRP problems across four scales: 100, 200, 500, and 1000. The dataset at the 100-scale contains 1,000 instances, while the others contain 128 instances each. In these settings, the vehicle capacity is uniformly set to 50, the duration limit is set to 3, the time window for the depot node in Time Window problems is [0, 4.6], and for backhaul problems, the proportion of backhaul nodes is set to 20%. Our datasets are generated based on the MVMoE codebase to evaluate the performance of various baselines.
Training Configuration and Model Parameters
We follow a similar setup to MVMoE, training our model on 6 VRPs and evaluating its zero-shot generalization performance on 10 unseen tasks (refer to Figure 1). The training process consists of two phases: teacher model pre-training and student model training.(1) Teacher Model Pre-training: We employ the POMO model as our teacher model. For each of the 6 visible tasks, we independently traine a single-task model using reinforcement learning. All teacher models share the same parameter configuration: 6 encoder layers, 1 decoder layer, an embedding dimension of 128, 8 attention heads, and a Feedforward hidden dimension of 512. The teacher models are trained for 4000 epochs on random instances of scale 100, with 20,000 training instances per epoch, a batch size of 128, and an initial learning rate of 0.0001. (2) Student Model Training: The student model consists of 1 encoder layer and 6 decoder layers, with embedding dimensions set to 128 and 96, employing 8 multi-head attention heads and a Feedforward hidden dimension of 512. The student model simultaneously learns from the 6 pre-trained teacher models on datasets with a problem scale of 100. The training batch size is set to 1500 (250 * 6), with 24,000 training instances per epoch, an initial learning rate of 0.0001, and a total of 900 training epochs. After the 300th epoch, the learning rate is halved every 100 epochs.
Inference and Metrics
We adopt the Average Objective Value and Gap as evaluation metrics, where smaller values indicate better performance. The Objective Value represents the total distance of the solution in the Vehicle Routing Problem, while the Gap evaluates the performance difference of each method compared to a traditional baseline method (such as HGS-PyVRP).
We test the results of a Single Trajectory Greedy Search (S.T.) on all methods. Furthermore, to compare with the results of other solvers exploring multiple trajectories under data augmentation (denoted as M×$\times$×aug8, where M is the number of nodes and aug8 represents 8 types of data augmentation[8]), our method employs the R3C strategy iterated 200 times under a single trajectory
6.1 Main Results
Table 1: Performance on seen VRPs across three scales. * represents the baseline used for gap calculation.
Method Pro. Problem Size Pro. Problem Size
n=100 n=500 n=1k n=100 n=500 n=1k
HGS-PyVRP CVRP 15.53(*) 62.07(*) 119.54(*) VRPTW 24.35(*) 90.61(*) 166.47(*)
OR-Tools 15.94(2.63%) 66.51(7.15%) 125.91(5.32%) 25.21(3.54%) 98.45(8.65%) 178.47(7.21%)
MT-POMO(S.T.) 16.25(4.64%) 70.60(13.74%) 146.92(22.90%) 26.66(10.40%) 122.61(35.50%) 247.44(66.47%)
MT-POMO(M+aug8) 15.79(1.69%) 67.99(9.54%) 136.62(14.28%) 25.61(5.18%) 115.43(27.39%) 229.82(38.06%)
MVMoE(S.T.) 16.30(5.00%) 77.44(24.77%) 191.07(59.83%) 26.88(10.44%) 122.78(35.26%) 267.33(60.59%)
MVMoE(M+aug8) 15.76(1.50%) 73.61(18.59%) 176.40(47.57%) 25.51(4.78%) 116.67(28.76%) 253.35(52.19%)
MTL-KD96 16.06(3.41%) 64.57(4.02%) 123.88(3.63%) 26.18(7.52%) 100.17(10.55%) 188.94(13.50%)
MTL-KD128 16.04 (3.30%) 64.61(4.09%) 124.44 (4.09%) 26.13(7.32%) 99.05(9.31%) 184.92(11.09%)
MTL-KD(R3C200)128 15.76(1.48%) 63.63(2.51%) 122.06 (2.10%) 25.31(3.93%) 96.43(6.42%) 181.85(9.24%)
HGS-PyVRP VRPL 15.58(*) 63.55(*) 122.68(*) VRPB - - -
OR-Tools 16.00(2.69%) 67.53(6.26%) 128.16(4.46%) 11.97(*) 47.76(*) 88.57(*)
MT-POMO(S.T.) 16.29(4.52%) 71.46(12.66%) 149.5521.90(%) 12.47(4.21%) 50.08(4.85%) 99.56(12.41%)
MT-POMO(M+aug8) 15.85(1.67%) 68.95(8.49%) 138.97(13.27%) 12.04(0.63%) 48.49(1.52%) 95.35(7.64%)
MVMoE(S.T.) 16.36(4.94%) 78.79(23.97%) 192.29(56.74%) 12.42(3.80%) 71.99(50.75%) 186.49(110.54%)
MVMoE(M+aug8) 15.81(1.45%) 74.67(17.50%) 177.89(45.00%) 12.01(0.31%) 66.26(38.73%) 167.26(88.84%)
MTL-KD96 16.14(3.58%) 65.45(2.99%) 126.11(2.79%) 12.39(3.51%) 46.12(-3.44%) 86.90(-1.89%)
MTL-KD128 16.12(3.45%) 65.48(3.03%) 126.66(3.25%) 12.38(3.41%) 45.99(-3.70%) 86.99(-1.78%)
MTL-KD(R3C200)128 15.82(1.50%) 64.5246 (1.53%) 124.59(1.56%) 12.02(0.43%) 44.58(-6.66%) 83.66(-5.55%)
HGS-PyVRP OVRP 9.71(*) 35.30(*) 66.10(*) OVRPTW 13.95(*) 48.15(*) 82.98(*)
OR-Tools 9.84(1.38%) 37.81(7.11%) 70.38(6.48%) 14.38(3.06%) 52.48(9.00%) 90.03(8.49%)
MT-POMO(S.T.) 10.66(9.83%) 44.62(26.41%) 92.78(40.36%) 15.56(11.53%) 71.63(48.76%) 147.84(78.16%)
MT-POMO(M+aug8) 10.17(4.74%) 41.93(18.78%) 85.27(29.00%) 14.85(6.43%) 66.74(38.61%) 135.99(63.87%)
MVMoE(S.T.) 10.77(10.94%) 49.35(39.81%) 135.30(104.70%) 15.70(12.54%) 82.81(71.99%) 215.21(159.34%)
MVMoE(M+aug8) 10.14(4.42%) 46.23(30.97%) 116.82(76.74%) 14.78(5.90%) 76.28(58.43%) 195.64(135.76%)
MTL-KD96 10.41(7.21%) 38.73(9.72%) 72.73(10.03%) 15.11(8.32%) 53.74(11.61%) 94.12(13.42%)
MTL-KD128 10.46(7.19%) 38.89(10.17%) 72.70(9.98%) 15.08 (8.07%) 53.89(11.91%) 93.96(13.22%)
MTL-KD(R3C200)128 10.05(3.53%) 37.7934 (7.07%) 71.40(8.03%) 14.53 (4.12%) 52.08(8.15%) 92.40(11.35%)
Performance on seen Tasks
We evaluat the performance of our model on the training tasks, and the results are shown in Table 1. The results demonstrate that our proposed MTL-KD model exhibits excellent performance on the seen training tasks, particularly showcasing a more pronounced advantage when dealing with larger-scale problems. This validates the effectiveness of our proposed knowledge distillation training framework and model architecture.
Performance on Unseen Tasks
The zero-shot generalization experimental results on 10 unseen tasks (as shown in Table 2) indicate that our MTL-KD model outperforms the compared models in most cases, demonstrating remarkable cross-task generalization ability, especially with a significant advantage on large-scale problems. This validates that multi-task knowledge distillation training endows the model with strong generalization capabilities across different VRP variants.
Table 2: Performance on unseen VRPs across three scales. * represents the baseline used for gap calculation.
Method Pro. Problem Size Pro. Problem Size
n=100 n=500 n=1k n=100 n=500 n=1k
HGS-PyVRP OVRPL 9.67(*) 34.70(*) 65.38(*) VRPLTW 24.44(*) 91.86(*) 174.79(*)
OR-Tools 9.79(1.25%) 37.09(6.89%) 69.64(6.51%) 25.65(4.96%) 99.50(8.31%) 186.38(6.63%)
MT-POMO(S.T.) 10.61(9.69%) 43.77(26.13%) 92.49(41.45%) 26.76(9.52%) 123.97(34.95%) 252.82(44.65%)
MT-POMO(M+aug8) 10.13(4.72%) 41.28(18.95%) 84.40(29.08%) 25.71(5.21%) 116.62(26.95%) 235.94(34.99%)
MVMoE(S.T.) 10.76(11.25%) 48.60(40.05%) 135.08(106.58%) 27.00(10.48%) 124.25(35.25%) 274.68(57.15%)
MVMoE(M+aug8) 10.10(4.43%) 45.58(31.36%) 116.71(78.48%) 25.62(4.84%) 118.00(28.45%) 260.07(48.79%)
MTL-KD(S.T.)96 10.42(7.73%) 38.23(10.17%) 72.32(10.61%) 26.37 (7.89%) 101.39(10.36%) 195.65(11.94%)
MTL-KD(S.T.)128 10.44(7.95%) 38.43(10.76%) 72.78(11.31%) 26.35(7.81%) 100.40 (9.30%) 192.41(10.08%)
MTL-KD(R3C200)128 10.08(4.29%) 37.27(7.41%) 71.35(9.12%) 25.55(4.54%) 97.85(6.51%) 189.02 (8.14%)
HGS-PyVRP OVRPLTW 14.00(*) 47.97 (*) 83.68(*) OVRPB - - -
OR-Tools 14.28(1.98%) 52.94(10.36%) 91.69(9.57%) 8.37(*) 29.98 (*) 54.87 (*)
MT-POMO(S.T.) 15.61(11.46%) 71.24(48.52%) 147.97(76.82%) 9.52(13.82%) 35.43(18.17%) 74.59(35.95%)
MT-POMO(M+aug8) 14.90(6.39%) 66.65(38.95%) 137.13(63.87%) 8.98(7.34%) 32.76(9.28%) 66.91(21.95%)
MVMoE(S.T.) 15.74(12.39%) 82.49(71.97%) 219.21(161.95%) 9.74(16.41%) 45.80(52.76%) 129.38(135.81%)
MVMoE(M+aug8) 14.83(5.90%) 76.12(58.68%) 197.84(136.41%) 8.96(7.10%) 40.41(34.77%) 109.08(98.81%)
MTL-KD(S.T.)96 15.28 (9.12%) 53.38 (11.28%) 94.65 (13.11%) 9.25(10.58%) 30.90(3.08%) 58.57(6.76%)
MTL-KD(S.T.)128 15.26(8.95%) 53.56(11.65%) 95.33(13.92%) 9.27(10.81%) 30.45 (1.58%) 56.92 (3.73%)
MTL-KD(R3C200)128 14.71(5.02%) 51.97(8.34%) 93.91(12.22%) 8.78(4.95%) 28.73(-4.16%) 52.41(-4.48%)
OR-Tools VRPBL 12.02(*) 47.93 (*) 89.82 (*) VRPBTW 25.41(*) 97.77(*) 194.69(*)
MT-POMO(S.T.) 12.71(5.79%) 50.70(5.77%) 99.51(10.78%) 28.64(12.70%) 126.00(28.87%) 270.68(39.03%)
MT-POMO(M+aug8) 12.10(0.66%) 48.14(0.45%) 94.26(4.94%) 26.94(6.04%) 118.20(20.89%) 251.58(29.22%)
MVMoE(S.T.) 12.86(7.06%) 70.03(46.11%) 159.91(78.03%) 28.88(13.67%) 125.89(28.76%) 294.23(51.13%)
MVMoE(M+aug8) 12.05(0.28%) 63.57(32.63%) 145.61(62.10%) 26.89(5.82%) 119.44(22.16%) 276.17(41.85%)
MTL-KD(S.T.)96 12.56(4.54%) 45.90(-4.23%) 86.45(-3.75%) 28.48(12.11%) 105.43(7.83%) 215.94(10.92%)
MTL-KD(S.T.)128 12.54(4.32%) 45.67(-4.72%) 86.28(-3.95%) 28.49 (12.12%) 103.85 (6.21%) 209.82(7.77%)
MTL-KD(R3C200)128 12.08(0.55%) 44.27(-7.64%) 82.56(-8.08%) 26.64(4.84%) 100.25(2.53%) 205.13(5.37%)
OR-Tools OVRPBL 8.35 (*) 29.60(*) 54.30 (*) OVRPBTW 14.38(*) 51.88 (*) 90.86 (*)
MT-POMO(S.T.) 9.50(13.79%) 35.34(19.37%) 74.83(37.80%) 16.92(17.64%) 72.91(40.53%) 152.63(67.98%)
MT-POMO(M+aug8) 8.96(7.35%) 32.60(10.13%) 66.87(23.13%) 15.88(10.39%) 67.95(30.96%) 140.72(54.87%)
MVMoE(S.T.) 9.75(16.75%) 45.46(53.57%) 127.34(134.49%) 17.07(18.68%) 81.57(57.22%) 214.76(136.36%)
MVMoE(M+aug8) 8.94(7.12%) 40.35(36.29%) 108.53(99.86%) 15.81(9.90%) 74.43(43.46%) 193.33(112.77%)
MTL-KD(S.T.)96 9.29(11.27%) 31.96(7.96%) 61.37(13.01%) 16.70(16.10%) 56.02(7.97%) 99.85(9.89%)
MTL-KD(S.T.)128 9.41(12.75%) 31.42(6.14%) 59.25(9.11%) 16.74(16.39%) 56.02(7.98%) 99.20(9.18%)
MTL-KD(R3C200)128 8.84(5.95%) 29.23(-1.26%) 53.77(-0.98%) 15.58(8.30%) 53.75(3.62%) 97.37(7.17%)
OR-Tools VRPBLTW 25.34 (*) 103.17(*) 189.31(*) OVRPBLTW 14.25(*) 52.37(*) 92.16(*)
MT-POMO(S.T.) 28.92(14.11%) 131.68(27.64%) 264.39(39.66%) 16.78(17.79%) 73.26(39.90%) 152.61(65.60%)
MT-POMO(M+aug8) 27.25(7.52%) 123.71(19.91%) 245.61(29.74%) 15.74(10.45%) 68.08(30.01%) 141.27(53.28%)
MVMoE(S.T.) 29.19(15.18%) 132.23(28.17%) 286.27(51.21%) 16.95(18.94%) 82.36(57.28%) 215.69(134.03%)
MVMoE(M+aug8) 27.14(7.10%) 125.27(21.43%) 269.13(42.16%) 15.67(9.97%) 74.83(42.90%) 195.45(112.07%)
MTL-KD(S.T.)96 29.01(14.4703%) 111.33(7.92%) 210.45(11.17%) 16.78(17.74%) 56.34(7.59%) 101.43 (10.06%)
MTL-KD(S.T.)128 29.13 (14.96%) 110.33 (6.94%) 205.52 (8.56%) 16.95(18.93%) 57.15(9.13%) 101.64(10.29%)
MTL-KD(R3C200)128 27.12(7.03%) 106.52(3.26%) 201.29(6.33%) 15.65 (9.84%) 54.68(4.41%) 99.64(8.11%)
Performance on Real-World Instances
We also evaluate the performance of our model on real-world datasets, including Set-X (medium and large scale) from CVRPLIB[25] for CVRP and the Solomon dataset[26] for VRPTW. We compare our model against several single-task and multi-task models, and the experimental results are presented in Tables 5, 6, and 7, AppendixC. Our model consistently outperforms other single-task and multi-task models across all three real-world datasets, demonstrating particularly strong performance on the large-scale Set-X dataset, where the gap is significantly smaller than other baselines. This indicates the strong applicability of our approach to real-world scenarios.
6.2 Ablation Study
Scale Generalization Comparison: Student vs. Teacher
While heavy decoders are generally considered to have significant potential for large-scale generalization, our model’s learning teacher model, POMO, primarily excels at its specific training scale and lacks cross-scale generalization capabilities. This naturally raises a crucial question: will the student model also be limited to learning knowledge specific to a particular scale, thereby restricting its generalization ability? To investigate this, we compare the performance of several single-task teacher models with our MTL-KD model on the same dataset, as shown in Figure 4. In contrast to the teacher models, the student model demonstrates a notable capacity for scale generalization. This observation indicates that the MTL-KD model, during the learning process, does not simply replicate the teacher’s knowledge but rather effectively adapts to variations in scale, consequently achieving robust generalization performance.
Figure 4: Performance Comparison between Teacher and Student Models, both trained on instances of scale 100.
Performance Comparison: KD vs. RL
To highlight the effectiveness of our knowledge distillation approach, we compared the MTL-KD model with a LEHD model trained using RL. Given the high computational cost associated with training the LEHD model with RL, we were only able to train it at a scale of 20. Subsequently, we evaluated the average gap of both models on all seen and unseen tasks, with the experimental results presented in Table 3. The results indicate that the RL method, due to its limitation to training on small scales, struggles to leverage the full potential of the LEHD model, leading to significantly poor performance. In contrast, our distillation-based MTL-KD model demonstrates a clear performance advantage.
Table 3: Heavy Decoder Performance (RL vs. KD) on Seen/Unseen Tasks (Average Gap).
Problem Size 100 200 500 1000
Training task RLMT-LEHD 21.7834% 27.2457% 38.2808% 50.0606%
MTL-KD 5.4569% 5.2775% 5.8028% 6.6413%
unseen task RLMT-LEHD 31.2513% 37.1404% 48.9074% 62.2226%
MTL-KD 19.1658% 12.9145% 10.8275% 13.3322%
Effectiveness Analysis of R3C
To validate the effectiveness of our proposed R3C method, we conduct experiments on CVRP, VRPL, OVRP, and VRPTW at a scale of 100. We perform an ablation study on the random reordering of the subtour external order component, with the following comparative experiments: random sampling on the original solution only (RS); random reordering of the subtour external order followed by sampling (RS+Ro); and for CVRP and VRPL, we also analyze the impact of adding a random flipping operation (F) since it does not affect the legality of the solution. The experimental results are shown in Figure 5. Incorporating the random reordering operation significantly improves the iterative performance during reconstruction, which benefits from the ability of random reordering to enhance the diversity of sampled subproblems. Furthermore, the random flipping operation can moderately increase the diversity of sampled solutions when only initial solution random sampling is performed; however, its impact is minimal with random reordering, further confirming the effectiveness of the random reordering operation.
Figure 5: Impact of Different Components in R3C.
7 Conclusion
This paper implements a heavy decoder model in the multi-task VRP domain and achieves label-free training of this heavy decoder model through a multi-task knowledge distillation method. The proposed approach demonstrates excellent performance on seen tasks, unseen tasks, and real-world datasets, exhibiting significant large-scale generalization ability, thereby validating the effectiveness of our method. Furthermore, the proposed R3C method further enhances the model’s performance. However, the structure of the heavy decoder leads to high computational complexity, which is a limitation. Future work will explore how to implement more efficient and robust model architectures in the multi-task domain.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflectthe contributions and scope of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses the limitations of the work in the last section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We will make our code, datasets, and pre-trained models publicly available upon final manuscript submission, with all implementation details thoroughly described in the paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We will release the dataset and code upon manuscript acceptance, with all experimental details documented in the paper to ensure full reproducibility of our results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Both the testing and training settings are detailed in Section Experiment 6
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Neural Combinatorial Optimization (NCO) methods typically only report the mean or the gap to the optimal solution (or the best-known solution). In this method, there is no randomness involved, and all random seeds are fixed at 2025.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe the required computational resources for the experiments in the experimental section 6.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This paper strictly adheres to the standards of academic writing, especially the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: This work focuses on solving and optimizing combinatorial optimization (CO) problems, which holds the potential to significantly enhance real-world applications. It appears to have no negative social impact. More discussion refers to Appendix E.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We list the licenses in Appendix D.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not introduce new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research involving human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing or research involving human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The paper does not describe the usage of LLMs as an important, original, or non-standard component of the core methods in this research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Related work
A.1 Autoregressive Neural Solvers
Autoregressive neural solvers have emerged as a prominent research direction in the field of NCO. These methods typically employ an encoder-decoder architecture, where the encoder generates node embeddings, and the decoder dynamically produces a sequence of nodes.
Early works were based on Pointer Networks[27], gradually evolving from supervised learning to reinforcement learning training[28, 29]. Subsequently, the Transformer framework was introduced and has become a mainstream paradigm[7]. POMO[8] further leverages the symmetry of the VRP, optimizing the baseline calculation in RL training and enhancing the exploration diversity of the model. Subsequent research has largely focused on improving the Transformer framework [30, 31, 10, 9, 32, 33, 34]. However, these approaches predominantly adopt a heavy encoder light decoder structure, which poses challenges in handling large-scale problems. To enhance the capability of addressing large instances, some methods continue to build upon HELD. Examples include ELG[35], which integrates local search strategies; INVIT[36], which utilizes K-nearest neighbor multi-view embeddings to enhance local information; and ICAM[37], which introduces instance-conditional adaptation to better perceive problem scales. On the other hand, some studies[15, 16] have explored heavy decoder architectures. By dynamically capturing the relationships between remaining nodes during the decoding process, these methods demonstrate superior generalization performance on large-scale problems. However, their high computational demands, often necessitating training on labeled data, represent a significant limitation.Furthermore, decomposition-based methods[38, 39, 40, 17] warrant attention. These approaches decompose large-scale problems into multiple smaller subproblems for solving, and autoregressive solvers can be employed at the lower level to address these subproblems.
A.2 Multi-Task Learning for Solving VRPs
To enhance the generalizability of neural solvers across diverse VRP variants, several existing methods have proposed unified multi-task models. For instance, Wang et al.[41] employs a multi-armed bandit approach to achieve efficient training across multiple combinatorial optimization problems. Lin et al.[42] pre-train a Transformer backbone on the Traveling Salesman Problem and fine-tune it for specific VRP variants, thereby extending its applicability to a subset of VRP problems. These methods represent initial attempts at cross-problem learning but are limited in scope, focusing primarily on a small number of problems. Building on these foundations, MT-POMO [11] views VRP variants as combinations of distinct attributes, enabling unified representation and zero-shot generalization across multiple VRP problems, achieving notable performance on ten tasks. MVMoE [12] further boosts model capacity and multi-task performance by introducing a mixture-of-experts model. RouteFinder[13] enhances multi-task training efficiency and performance through mixed batch training and enables rapid adaptation to unseen tasks via efficient adapter layers. CaDA [43] improves the model’s ability to handle diverse tasks within a multi-task framework by incorporating constraint awareness and a dual-branch structure. However, these studies have predominantly focused on small-scale problems, lacking extensive exploration of large-scale generalization capabilities.
Appendix B Problem Definition and Settings
Our VRP variant problem settings follow the work of Zhou et al. [12]. Detailed settings are as follows:
  • Coordinates: All node coordinates are uniformly sampled from the 2D space $[0,1)\times[0,1)$.
  • Demands: For all customer nodes, demands are randomly sampled from the set $\{1,2,\dots,9\}$.
  • Vehicle Capacity: Each vehicle has a fixed transportation capacity, which we set to 50 across all problem scales. The sum of demands for nodes visited by a vehicle must be less than or equal to its current capacity.
  • Open Routes: In open vehicle routing problems, vehicles are not required to return to the depot. We use a binary flag vector to distinguish whether a path is open; setting it to 1 indicates an open path.
  • Backhauls: In problems involving backhauls, a portion of the nodes have negative demands, referred to as backhaul nodes. We designate 20% of the customer nodes as backhaul nodes, whose demands are randomly sampled from $\{-1,-2,\dots,-9\}$.
  • Duration Limit: Each vehicle has a maximum travel distance limit, which we set to 3.
  • Time Windows: Customer node time windows are uniformly sampled, and service time is uniformly set to 0.2. Vehicles must visit nodes within their time windows; if a vehicle arrives early, it must wait until the earliest start time. The depot’s time window is set to $[0,3]$, with a service time of 0. Vehicle speed is uniformly set to 1, meaning the time spent traveling on a path equals its distance divided by speed. Notably, for non-open problems, it is crucial to ensure that after visiting the last customer node, the vehicle can return to the depot within its latest allowable time window; this consideration is not necessary for open problems.
By combining different constraints, 16 VRP variants can be formed, as detailed in Table 4.
Table 4: 16 VRP variants with five constraints.
Constraint Capacity Open Route Backhaul Duration Limit Time Window
CVRP
OVRP
VRPB
VRPL
VRPTW
OVRPTW
OVRPB
OVRPL
VRPBL
VRPBTW
VRPLTW
OVRPBL
OVRPBTW
OVRPLTW
VRPBLTW
OVRPBLTW
Appendix C Performance on Real-World Instances
Our test results on Real-World Instances datasets are presented in Tables 5, 6, and 7. All models are trained on instances of size $N=100$. During inference, data augmentation is applied, and a greedy rollout strategy is adopted by default. Some results for comparative models are sourced from MVMoE [12] and RouteFinder [13].
Table 5: Results on small-scale CVRPLIB instances (Set-X) [25]. All models are trained on instances of size $N=100$, and inference utilized data augmentation[8] and greedy rollout by default.
Set-X POMO POMO-MTL MVMoE/4E MVMOE/4E-L MTL-KD
Instance Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap
X-n101-k25 27591 30138 9.231% 32482 17.727% 29361 6.415% 29015 5.161% 29058 5.317%
X-n106-k14 26362 39322 49.162% 27369 3.820% 27278 3.475% 27242 3.338% 26918 2.109%
X-n110-k13 14971 15223 1.683% 15151 1.202% 15089 0.788% 15196 1.503% 15407 2.912%
X-n115-k10 12747 16113 26.406% 14785 15.988% 13847 8.629% 13325 4.534% 13513 6.009%
X-n120-k6 13332 14085 5.648% 13931 4.493% 14089 5.678% 13833 3.758% 13657 2.438%
X-n125-k30 55539 58513 5.355% 60687 9.269% 58944 6.131% 58603 5.517% 57615 3.738%
X-n129-k18 28940 29246 1.057% 30332 4.810% 29802 2.979% 29457 1.786% 29309 1.275%
X-n134-k13 10916 11302 3.536% 11581 6.092% 11353 4.003% 11398 4.416% 11363 4.095%
X-n139-k10 13590 14035 3.274% 13911 2.362% 13825 1.729% 13800 1.545% 13911 2.362%
X-n143-k7 15700 16131 2.745% 16660 6.115% 16125 2.707% 16147 2.847% 15955 1.624%
X-n148-k46 43448 49328 13.533% 50782 16.880% 46758 7.618% 45599 4.951% 45463 4.638%
X-n153-k22 21220 32476 53.040% 26237 23.643% 23793 12.125% 23316 9.877% 23340 9.991%
X-n157-k13 16876 17660 4.646% 17510 3.757% 17650 4.586% 17410 3.164% 17161 1.689%
X-n162-k11 14138 14889 5.312% 14720 4.117% 14654 3.650% 14662 3.706% 14487 2.469%
X-n167-k10 20557 21822 6.154% 21399 4.096% 21340 3.809% 21275 3.493% 21053 2.413%
X-n172-k51 45607 49556 8.659% 56385 23.632% 51292 12.465% 49073 7.600% 47850 4.918%
X-n176-k26 47812 54197 13.354% 57637 20.549% 55520 16.121% 52727 10.280% 52476 9.755%
X-n181-k23 25569 37311 45.923% 26219 2.542% 26258 2.695% 26241 2.628% 25919 1.369%
X-n186-k15 24145 25222 4.461% 25000 3.541% 25182 4.295% 24836 2.862% 24711 2.344%
X-n190-k8 16980 18315 7.862% 18113 6.673% 18327 7.933% 18113 6.673% 17539 3.292%
X-n195-k51 44225 49158 11.154% 54090 22.306% 49984 13.022% 48185 8.954% 46301 4.694%
X-n200-k36 58578 64618 10.311% 61654 5.251% 61530 5.039% 61483 4.959% 60978 4.097%
X-n209-k16 30656 32212 5.076% 32011 4.420% 32033 4.492% 32055 4.564% 31536 2.871%
X-n219-k73 117595 133545 13.564% 119887 1.949% 121046 2.935% 120421 2.403% 118499 0.769%
X-n228-k23 25742 48689 89.142% 33091 28.549% 31054 20.636% 28561 10.951% 28156 9.378%
X-n237-k14 27042 29893 10.543% 28472 5.288% 28550 5.577% 28486 5.340% 27789 2.762%
X-n247-k50 37274 56167 50.687% 45065 20.902% 43673 17.167% 41800 12.143% 41106 10.281%
X-n251-k28 38684 40263 4.082% 40614 4.989% 41022 6.044% 40822 5.527% 39877 3.084%
Avg. Gap 16.629% 9.820% 6.884% 5.160% 4.025%
Table 6: Results on VRPTW instances (Set-Solomon) [26]. All models are trained on instances of size $N=100$, and inference utilized data augmentation[8] and greedy rollout by default.
Set-Solomon POMO POMO-MTL MVMoE/4E MVMOE/4E-L MTL-KD
Instance Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap
R101 1637.7 1805.6 10.252% 1821.2 11.205% 1798.1 9.794% 1730.1 5.641% 1768.0 7.956%
R102 1466.6 1556.7 6.143% 1596.0 8.823% 1572.0 7.187% 1574.3 7.345% 1564.4 6.668%
R103 1208.7 1341.4 10.979% 1327.3 9.812% 1328.2 9.887% 1359.4 12.470% 1379.0 14.090%
R104 971.5 1118.6 15.142% 1120.7 15.358% 1124.8 15.780% 1098.8 13.100% 1074.1 10.561%
R105 1355.3 1506.4 11.149% 1514.6 11.754% 1479.4 9.157% 1456.0 7.433% 1465.2 8.109%
R106 1234.6 1365.2 10.578% 1380.5 11.818% 1362.4 10.352% 1353.5 9.627% 1346.1 9.031%
R107 1064.6 1214.2 14.052% 1209.3 13.592% 1182.1 11.037% 1196.5 12.391% 1193.9 12.145%
R108 932.1 1058.9 13.604% 1061.8 13.915% 1023.2 9.774% 1039.1 11.481% 1035.4 11.082%
R109 1146.9 1249.0 8.902% 1265.7 10.358% 1255.6 9.478% 1224.3 6.750% 1260.0 9.861%
R110 1068.0 1180.4 10.524% 1171.4 9.682% 1185.7 11.021% 1160.2 8.635% 1199.5 12.313%
R111 1048.7 1177.2 12.253% 1211.5 15.524% 1176.1 12.148% 1197.8 14.220% 1178.7 12.396%
R112 948.6 1063.1 12.070% 1057.0 11.427% 1045.2 10.183% 1044.2 10.082% 1057.2 11.448%
RC101 1619.8 2643.0 63.168% 1833.3 13.181% 1774.4 9.544% 1749.2 7.988% 1774.0 9.520%
RC102 1457.4 1534.8 5.311% 1546.1 6.086% 1544.5 5.976% 1556.1 6.771% 1588.4 8.989%
RC103 1258.0 1407.5 11.884% 1396.2 10.986% 1402.5 11.486% 1415.3 12.502% 1451.8 15.405%
RC104 1132.3 1261.8 11.437% 1271.7 12.311% 1265.4 11.755% 1264.2 11.649% 1241.6 9.653%
RC105 1513.7 1612.9 6.553% 1644.9 8.668% 1635.5 8.047% 1619.4 6.980% 1688.3 11.535%
RC106 1372.7 1539.3 12.137% 1552.8 13.120% 1505.0 9.638% 1509.5 9.968% 1468.8 7.001%
RC107 1207.8 1347.7 11.583% 1384.8 14.655% 1351.6 11.906% 1324.1 9.625% 1367.6 13.231%
RC108 1114.2 1305.5 17.169% 1274.4 14.378% 1254.2 12.565% 1247.2 11.939% 1232.3 10.600%
RC201 1261.8 2045.6 62.118% 1761.1 39.570% 1577.3 25.004% 1517.8 20.285% 1490.8 18.149%
RC202 1092.3 1805.1 65.257% 1486.2 36.062% 1616.5 47.990% 1480.3 35.520% 1336.5 22.356%
RC203 923.7 1470.4 59.186% 1360.4 47.277% 1473.5 59.521% 1479.6 60.182% 1237.8 34.005%
RC204 783.5 1323.9 68.973% 1331.7 69.968% 1286.6 64.212% 1232.8 57.342% 1066.0 36.056%
RC205 1154.0 1568.4 35.910% 1539.2 33.380% 1537.7 33.250% 1440.8 24.850% 1448.3 25.503%
RC206 1051.1 1707.5 62.449% 1472.6 40.101% 1468.9 39.749% 1394.5 32.671% 1354.7 28.884%
RC207 962.9 1567.2 62.758% 1375.7 42.870% 1442.0 49.756% 1346.4 39.831% 1235.9 28.352%
RC208 776.1 1505.4 93.970% 1185.6 52.764% 1107.4 42.688% 1167.5 50.437% 1064.0 37.096%
Avg. Gap 28.054% 21.380% 20.317% 18.490% 15.786%
Table 7: Results on large-scale CVRPLIB instances (Set-X) [25]. All models are trained on instances of size $N=100$, and inference utilized data augmentation[8] and greedy rollout by default.
Set-X POMO LEHD POMO-MTL MVMoE/4E MVMOE/4E-L RF-MVMOE RF-TE MTL-KD
Instance Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap
X-n502-k39 69226 75617 9.232% 71438 3.195% 77284 11.640% 73533 6.222% 74429 7.516% 76338 10.274% 71791 3.705% 71124 2.742%
X-n513-k21 24201 30518 26.102% 25624 5.880% 28510 17.805% 32102 32.647% 31231 29.048% 32639 34.866% 28465 17.619% 25947 7.215%
X-n524-k153 154593 201877 30.586% 280556 81.480% 192249 24.358% 186540 20.665% 182392 17.982% 170999 10.612% 174381 12.800% 171306 10.811%
X-n536-k96 94846 106073 11.837% 103785 9.425% 106514 12.302% 109581 15.536% 108543 14.441% 105847 11.599% 103272 8.884% 101893 7.430%
X-n548-k50 86700 103093 18.908% 90644 4.549% 94562 9.068% 95894 10.604% 95917 10.631% 104289 20.287% 100956 16.443% 89169 2.848%
X-n561-k42 42717 49370 15.575% 44728 4.708% 47846 12.007% 56008 31.114% 51810 21.287% 53383 24.969% 49454 15.771% 45467 6.438%
X-n573-k30 50673 83545 64.871% 53482 5.543% 60913 20.208% 59473 17.366% 57042 12.569% 61524 21.414% 55952 10.418% 53466 5.512%
X-n586-k159 190316 229887 20.792% 232867 22.358% 208893 9.761% 215668 13.321% 214577 12.748% 212151 11.473% 205575 8.018% 200863 5.542%
X-n599-k92 108451 150572 38.839% 115377 6.386% 120333 10.956% 128949 18.901% 125279 15.517% 126578 16.714% 116560 7.477% 113513 4.668%
X-n613-k62 59535 68451 14.976% 62484 4.953% 67984 14.192% 82586 38.718% 74945 25.884% 73456 23.383% 67267 12.987% 63035 5.879%
X-n627-k43 62164 84434 35.825% 67568 8.693% 73060 17.528% 70987 14.193% 70905 14.061% 70414 13.271% 67572 8.700% 65755 5.777%
X-n641-k35 63682 75573 18.672% 68249 7.172% 72643 14.071% 75329 18.289% 72655 14.090% 71975 13.023% 70831 11.226% 67593 6.141%
X-n655-k131 106780 127211 19.134% 117532 10.069% 116988 9.560% 117678 10.206% 118475 10.952% 119057 11.497% 112202 5.078% 109748 2.780%
X-n670-k130 146332 208079 42.197% 220927 50.977% 190118 29.922% 197695 35.100% 183447 25.364% 168226 14.962% 168999 15.490% 161076 10.076%
X-n685-k75 68205 79482 16.534% 72946 6.951% 80892 18.601% 97388 42.787% 89441 31.136% 82269 20.620% 77847 14.137% 73249 7.395%
X-n701-k44 81923 97843 19.433% 86327 5.376% 92075 12.392% 98469 20.197% 94924 15.870% 90189 10.090% 89932 9.776% 85967 4.936%
X-n716-k35 43373 51381 18.463% 46502 7.214% 52709 21.525% 56773 30.895% 52305 20.593% 52250 20.467% 49669 14.516% 47012 8.390%
X-n733-k159 136187 159098 16.823% 149115 9.493% 161961 18.925% 178322 30.939% 167477 22.976% 156387 14.833% 148463 9.014% 142712 4.791%
X-n749-k98 77269 87786 13.611% 83439 7.985% 90582 17.229% 100438 29.985% 94497 22.296% 92147 19.255% 85171 10.227% 82295 6.505%
X-n766-k71 114417 135464 18.395% 131487 14.919% 144041 25.891% 152352 33.155% 136255 19.086% 130505 14.061% 129935 13.563% 123310 7.772%
X-n783-k48 72386 90289 24.733% 76766 6.051% 83169 14.897% 100383 38.677% 92960 28.423% 96336 33.087% 83185 14.919% 77332 6.833%
X-n801-k40 73305 124278 69.536% 77546 5.785% 85077 16.059% 91560 24.903% 87662 19.585% 87118 18.843% 86164 17.542% 78041 6.460%
X-n819-k171 158121 193451 22.344% 178558 12.925% 177157 12.039% 183599 16.113% 185832 17.525% 179596 13.581% 174441 10.321% 170672 7.938%
X-n837-k142 193737 237884 22.787% 207709 7.212% 214207 10.566% 229526 18.473% 221286 14.220% 230362 18.904% 208528 7.635% 203165 4.866%
X-n856-k95 88965 152528 71.447% 92936 4.464% 101774 14.398% 99129 11.425% 106816 20.065% 105801 18.924% 98291 10.483% 94079 5.748%
X-n876-k59 99299 119764 20.609% 104183 4.918% 116617 17.440% 119619 20.463% 114333 15.140% 114016 14.821% 107416 8.174% 105873 6.620%
X-n895-k37 53860 70245 30.421% 58028 7.739% 65587 21.773% 79018 46.710% 64310 19.402% 69099 28.294% 64871 20.444% 58606 8.812%
X-n916-k207 329179 399372 21.324% 385208 17.021% 361719 9.885% 383681 16.557% 374016 13.621% 373600 13.494% 352998 7.236% 346940 5.396%
X-n936-k151 132715 237625 79.049% 196547 48.097% 186262 40.347% 220926 66.466% 190407 43.471% 161343 21.571% 163162 22.942% 152094 14.602%
X-n957-k87 85465 130850 53.104% 90295 5.651% 98198 14.898% 113882 33.250% 105629 23.593% 123633 44.659% 102689 20.153% 90407 5.782%
X-n979-k58 118976 147687 24.132% 127972 7.561% 138092 16.067% 146347 23.005% 139682 17.404% 131754 10.740% 129952 9.225% 127650 7.291%
X-n1001-k43 72355 100399 38.759% 76689 5.990% 87660 21.153% 114448 58.176% 94734 30.929% 88969 22.962% 85929 18.760% 78833 8.953%
Avg. Gap 29.658% 12.836% 16.796% 26.408% 19.607% 18.795% 12.303% 6.655%
Appendix D Licenses for Code and Datasets
The licenses for the codes and the datasets used in this work are listed in Table 8.
Table 8: Licenses of Used Codes and Datasets
Type Resource License Type URL / Reference
Code MVMoE[12] MIT License https://github.com/RoyalSkye/Routing-MVMoE
Code MT-POMO[11] MIT License https://github.com/FeiLiu36/MTNCO
Code HGS-PyVRP[24] MIT License https://github.com/PyVRP/PyVRP
Code OR-Tools[6] Apache License https://github.com/google/or-tools
Dataset CVRPLIB(Set-X)[25] Available for academic research use http://vrp.galgos.inf.puc-rio.br/index.php/en/
Dataset Solomon[26] Available for academic research use https://www.sintef.no/projectweb/top/vrptw/solomon-benchmark/
Appendix E Broader Impacts
This research contributes to the field of neural combinatorial optimization by employing advanced machine learning techniques to address large-scale and multi-variant VRP problems. We believe that the proposed multi-task knowledge distillation training framework, heavy decoder model, and Random Reordering Reconstruction (R3C) strategy can provide valuable insights and inspire subsequent work to explore more efficient and effective neural methods for solving large-scale and diverse VRP problems. As a general learning-based approach to solving the VRP, the proposed multi-task knowledge distillation training framework and R3C strategy do not inherently possess any specific potential negative social impacts.